# Channel Knowledge Map Construction via Physics-Inspired Diffusion Model Without Prior Observations

### v1

@@ Abstract
The ability to construct Channel Knowledge Map (CKM) with high precision is essential for environment awareness in 6G wireless systems. However, most existing CKM construction methods formulate the task as an image super-resolution or generation problem, thereby employing models originally developed for computer vision. As a result, the generated CKMs often fail to capture the underlying physical characteristics of wireless propagation. In this paper, we focus on the construction of CKM for large-scale fading scenarios and design three physics-based constraint terms to characterize the spatial distribution patterns of large-scale fading. By integrating these physical constraints with a state-of-the-art diffusion model that possesses superior generative capability, a physics-inspired diffusion model for CKM construction is proposed. Following this motivation, we derive the loss function of the diffusion model augmented with physics-based constraint terms and further design the training and generation framework for the proposed physics-inspired CKM generation diffusion model. Extensive experiments show that our approach outperforms all existing methods in terms of construction accuracy. Moreover, the proposed model provides a unified and effective framework with strong potential for generating diverse, accurate, and physically consistent CKM.
I Introduction
@@
In sixth-generation (6G) communication systems, the concept of sensing is expected to evolve from spectrum-level sensing in cognitive radio to environment-level sensing in integrated sensing and communication (ISAC) systems, and further to semantic-level sensing for intelligent network resource optimization. Since electromagnetic waves experience attenuation, reflection, scattering, and diffusion during propagation, storing channel-related information at each spatial location enables the establishment of a link between the physical environment and the wireless communication system. Such spatially distributed data, referred to as the Channel Knowledge Map (CKM)[1], serves as a knowledge-driven representation of the propagation environment. CKM not only facilitates a deeper understanding of environmental information by communication systems but also provides essential support for semantic-level perception required in intelligent network resource management, thereby attracting significant attention in 6G research.
@@
CKMs can be categorized into various types according to the channel-related information they represent. Such information may include large-scale fading features such as line-of-sight (LOS) and non-line-of-sight (NLOS) states, path loss, and shadow fading, as well as small-scale fading features such as multipath delay, Doppler power spectrum, and angular power spectrum. In this work, we investigate the Channel Gain Map (CGM), a specific form of CKM that characterizes the path loss value at each spatial position within a region. In some studies, this type of map is also referred to as Radio Map. A more comprehensive discussion of the concepts of Radio Map and CKM can be found in surveys[2] and[3]. After obtaining the CKM, its major applications in wireless communication systems can be categorized into four types, as illustrated in Fig. 1: base station (BS) placement or deployment, unmanned aerial vehicle (UAV) trajectory planning, user resource scheduling, and indoor localization.
@@
Indoor positioning is one of the most prevalent application areas of CKM, as location information can be inferred from CKM based on the user’s online estimation of channel-related features. The authors in[4] and[5] respectively propose constructing a received signal strength (RSS) map and an angle–delay domain channel impulse response map for indoor localization. In addition, UAV path planning constitutes another important application area of CKM. As demonstrated in[6], LOS/NLOS state maps can be exploited to jointly optimize UAV trajectories and user scheduling, thereby maximizing the aggregated data throughput of all user nodes within a limited flight time. Furthermore,[7] shows that, by constructing a signal-to-interference-plus-noise ratio (SINR) map, the Dijkstra algorithm can be employed to identify the shortest path that consistently satisfies communication quality requirements while avoiding co-channel interference from neighboring terrestrial base stations. Meanwhile, existing studies have demonstrated that employing CKM in BS placement or configuration can yield notable performance gains and cost reductions. For example,[8] presents a CGM-based scheme that optimizes the antenna downtilt and azimuth angles to maximize cell coverage and [9] proposes a CGM-enabled base-station deployment method that satisfies multi-BS coverage requirements and quality-of-service constraints. Since CKMs can accurately characterize the spatial properties of wireless channels, their applications can be extended to higher-layer network functionalities such as user resource allocation and scheduling. The authors in[10] develop a neural-network–based power control scheme leveraging large-scale fading coefficients. Hence, constructing multi-cell CGMs can provide valuable support for resource allocation and power control in multi-cell massive MIMO systems. Moreover, as shown in[11], CKM can be employed in cell-free massive MIMO (CF-mMIMO) systems to facilitate access-point (AP) switching, thereby mitigating pilot contamination and reducing energy consumption associated with path-loss estimation.
In addition, CKMs can also play an important role in various other areas such as physical layer security[12], intelligent beamforming[13], and device-to-device communications[14].
@@
Considering that obtaining complete channel information over an entire spatial region purely through measurements is costly, time-consuming, and practically infeasible, research efforts have been devoted to developing efficient CKM construction methods. Depending on the implementation approach, existing CKM construction techniques can be categorized into four types: model-based methods using statistical channel models, model-based methods using deterministic channel models, spatial interpolation–based methods, and deep learning–based methods. The statistical channel model-based construction methods refer to utilizing existing statistical channel models to compute large-scale or small-scale fading channel information. Commonly used statistical channel models include the log-distance model[15] and the Okumura–Hata model[16] for large-scale fading computation, as well as Quadriga[17] and the 6G PCM (pervasive channel model)[18] for small-scale fading modeling. Although this method is simple, it often yields CKMs with limited accuracy. The deterministic channel model–based construction methods obtain channel-related information by performing ray-tracing–based channel simulations. Representative ray-tracing simulators include Wireless InSite[19], WinProp[20], and Sionna[21], among others. Although this approach can achieve high accuracy, it suffers from extremely high computational complexity, making it difficult to meet the real-time requirements of physical-layer algorithms.
@@
The spatial interpolation–based approach constructs CKMs using a limited number of precisely measured channel samples at selected locations. By exploiting the spatial correlation, smoothness, or low-rank characteristics of the wireless channel over the spatial domain, classical interpolation algorithms are applied to estimate the channel information in unmeasured regions. Typical spatial interpolation algorithms used for CKM construction include inverse distance weighting (IDW)[22], Kriging[23], radial basis function (RBF) interpolation[24], matrix completion[25], compressed sensing[26], and the expectation–maximization (EM) algorithm[27]. Although the above methods significantly reduce the computational complexity compared with ray-tracing–based approaches, they do not incorporate physical environmental information, and thus the resulting CKMs may not accurately reflect the actual propagation conditions. Moreover, these spatial interpolation algorithms rely on the assumptions of strong spatial correlation, smoothness, or low-rank characteristics of the channel. In densely built urban areas, where such assumptions are often violated, the accuracy of CKMs obtained through interpolation can degrade considerably.
@@
Deep learning models can effectively establish a mapping from the physical environment map to the CKM, thereby enabling more accurate and generalizable CKM construction. The earliest attempts to apply deep learning for CKM construction were inspired by models in the field of computer vision. Specifically, convolutional neural networks based on autoencoder architectures [28, 29, 30], UNet structures[31], and Laplacian pyramid frameworks[32] were introduced to capture the spatial correlations of large-scale fading channels. In addition, inspired by the outstanding performance of Transformer models in the field of computer vision, recent studies have introduced Transformer into CKM construction. As shown in [33] and[34], the self-attention mechanism of the Transformer can effectively capture long-range spatial correlations and overcome the limited global receptive field of convolutional neural networks, thereby achieving higher prediction accuracy.
@@
The discriminative models employed in the aforementioned studies typically yield over-smoothed predictions that lack sufficient high-frequency details. However, the spatial distribution of channel-related features in complex environments exhibits strong nonlinearity and fine-grained variations, making it difficult for purely supervised learning to accurately capture such characteristics. In contrast, generative models learn the underlying data distribution and are therefore capable of capturing higher-order statistical characteristics, enabling more accurate prediction of fine-grained spatial variations in channel information. Several studies[35, 36, 37] have developed conditional GAN (cGAN)–based frameworks to achieve environment-consistent CKM generation. However, the GAN framework involves a minimax optimization objective, which often leads to unstable training and convergence difficulties. In recent years, diffusion models have been recognized as a more powerful class of generative models, providing significantly higher generation quality than GANs. Inspired by their success in other domains, researchers have introduced conditional diffusion frameworks for CKM construction. The studies in[38, 39, 40] have demonstrated that conditional diffusion models can achieve state-of-the-art CKM prediction accuracy under various input conditions.
@@
However, the aforementioned studies that employ diffusion models for CKM construction directly transfer conditional diffusion frameworks from the field of computer vision. In image generation tasks, various improved conditional diffusion models are designed to produce images that are more realistic and visually appealing. Consequently, their performance is typically evaluated using image quality metrics such as the structural similarity index, Fréchet inception distance, and learned perceptual image patch similarity. In contrast, the primary objective in CKM construction is to achieve high prediction accuracy rather than visual fidelity. Therefore, conditional diffusion models with high image generation quality are not suitable for direct application to CKM construction. In this paper, motivated by the physical characteristics of wireless signal propagation, we integrate physical prior knowledge with the diffusion model and propose a physics-inspired diffusion-based CKM generation framework. The proposed model not only inherits the advantages of diffusion models in generating high-quality results with richer high-frequency details, but also ensures that the generated CKMs conform to the physical constraints of wireless signal propagation. It is worth noting that several physics-inspired methods for constructing CGM have already been proposed in[41] and[42]. These approaches introduce physical constraints derived from Maxwell’s equations and incorporate them into the neural network training process as additional loss terms. However, we note that Maxwell’s equations represent the microscopic formulation of wireless signal propagation, as the described electric field variations occur at the wavelength scale. Therefore, they are not well-suited to serve as physical constraints for large-scale fading, which characterizes the macroscopic propagation behavior of wireless signals. Moreover, the boundary conditions of electromagnetic reflections, refractions, and diffractions are highly complex, and existing physics-inspired methods do not impose boundary-condition residuals as training constraints, further reducing the physical fidelity of the CGM construction. To address these limitations, this paper heuristically designs three novel physical constraint terms that better characterize the spatial distribution of large-scale fading, and innovatively integrates them into the diffusion model framework.
@@
Based on the above discussion, the main contributions of this paper can be summarized as follows:
@@
    To address the insufficient physical consistency exhibited by existing CGM construction methods, we heuristically design three physics-inspired regularization terms, namely the edge loss, the regional propagation loss, and the multi-scale feature loss. These three regularizations respectively reflect distinct physical phenomena exhibited by electromagnetic waves when encountering obstacles, and are therefore closely tied to the propagation environment. This makes them particularly suitable for CGM construction in complex scenarios such as urban canyons and indoor environments.
@@
II CKM Construction Problem Description
The objective of this paper is to construct a classical type of CKM—the CGM. In the construction of CKM, including the CGM, the map is typically divided into an $H\times H$ grid. Consequently, both the ground-truth CGM $\mathbf{x}_{GT}$ and the predicted CGM $\mathbf{x}_{pred}$ are represented as real-valued matrices of size $H\times H$. Accordingly, when external conditional information is required for CGM construction, the input data are also represented as real-valued matrices of the same size. In the transmitter location matrix $\mathbf{L}_{\text{TX}}\in R^{H\times H}$, only the element corresponding to the transmitter position is set to 1, while all other elements are 0. In the building environment matrix $\mathbf{B}\in R^{H\times H}$, the elements representing building regions are set to 1, and all others are 0. In the vehicle location matrix $\mathbf{C}\in R^{H\times H}$, the element corresponding to the vehicle position is set to 1, with all remaining elements being 0. In the sparse observation matrix $\mathbf{S}\in R^{H\times H}$, the elements at the observation positions contain the actual observed values, whereas all other elements are set to 0.
@@ II CKM Construction Problem Description
III Theoretical Foundations of Diffusion Models
@@
III-A Denoising Diffusion Probabilistic Model
Inspired by nonequilibrium thermodynamics, Denoising Diffusion Probabilistic Model (DDPM)[43] formulates the forward noise-adding and reverse denoising processes as Markov chains. Therefore, DDPM mainly consists of two components: a forward process that gradually adds noise to the input, and a reverse process that learns how to denoise it.
@@ III-A Denoising Diffusion Probabilistic Model
$q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{x}_{0})$ $=\frac{q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1},\mathbf{x}_{0})q(\mathbf{x}_{t-1}\mid\mathbf{x}_{0})}{q(\mathbf{x}_{t}\mid\mathbf{x}_{0})}$ (6)
$\propto\mathcal{N}\left(\mathbf{x}_{t-1};\tilde{\mu}_{t},\tilde{\beta}_{t}\mathbf{I}\right),$
@@
III-B Denoising Diffusion Implicit Model
The DDPM requires $T$ iterative steps to generate a single sample, which results in relatively slow generation speed. To address this limitation, the Denoising Diffusion Implicit Model (DDIM)[44] removes the constraint that the reverse denoising process must follow a Markov chain, thereby achieving significantly faster sampling while maintaining comparable generation quality.

### v2

@@ Abstract
The ability to construct channel knowledge map (CKM) with high precision is essential for environment awareness in 6G wireless systems. However, most existing CKM construction methods formulate the task as an image super-resolution or generation problem, thereby employing models originally developed for computer vision. As a result, the generated CKMs often fail to capture the underlying physical characteristics of wireless propagation. In this paper, considering that acquiring channel observations incurs non-negligible time and cost, we focus on constructing CKM for large-scale fading scenarios without relying on prior observations, and we design three physics-based constraints to characterize the spatial distribution patterns of large-scale fading. By integrating these physical constraints with state-of-the-art diffusion model that possesses superior generative capability, a physics-inspired diffusion model for CKM construction is proposed. Following this motivation, we derive the loss function of the diffusion model augmented with physics-based constraint terms and further design the training and generation framework for the proposed physics-inspired CKM generation diffusion model. Extensive experiments show that our approach outperforms all existing methods in terms of construction accuracy. Moreover, the proposed model provides a unified and effective framework with strong potential for generating diverse, accurate, and physically consistent CKM.
Keywords:Channel knowledge map, diffusion model, large-scale fading, physics-inspired machine learning.
1 Introduction
@@
In sixth-generation (6G) communication systems, the concept of sensing is expected to evolve from spectrum-level sensing in cognitive radio to environment-level sensing in integrated sensing and communication (ISAC) systems, and further to semantic-level sensing for intelligent network resource optimization. Since electromagnetic waves experience attenuation, reflection, scattering, and diffusion during propagation, storing channel-related information at each spatial location enables the establishment of a link between the physical environment and the wireless communication system. Such spatially distributed data, referred to as channel knowledge map (CKM)[Zeng2021], serves as a knowledge-driven representation of the propagation environment. CKM not only facilitates a deeper understanding of environmental information by communication systems but also provides essential support for semantic-level perception required in intelligent network resource management, thereby attracting significant attention in 6G research.
@@
CKMs can be categorized into various types according to the channel-related information they represent. Such information may include large-scale fading features such as line-of-sight (LOS) and non-line-of-sight (NLOS) states, path loss, and shadow fading, as well as small-scale fading features such as multipath delay, Doppler power spectrum, and angular power spectrum. In this work, we investigate the channel gain map (CGM), a specific form of CKM that characterizes the path loss value at each spatial position within a region. In some studies, this type of map is also referred to as Radio Map. A more comprehensive discussion of the concepts of Radio Map and CKM can be found in surveys[Romero2022] and[Zeng2024]. After obtaining the CKM, its major applications in wireless communication systems can be categorized into four types, as illustrated in Fig. 1: base station (BS) placement or deployment, unmanned aerial vehicle (UAV) trajectory planning, user resource allocation, and radio-based localization.
@@
Radio-based localization is one of the most prevalent application areas of CKM, as location information can be inferred from CKM based on the user’s online estimation of channel-related features. The authors in[9031749] and[10025691] respectively propose constructing a received signal strength (RSS) map and an angle–delay domain channel impulse response map for localization. In addition, UAV path planning constitutes another important application area of CKM. As demonstrated in[8525324], LOS/NLOS state maps can be exploited to jointly optimize UAV trajectories and user scheduling, thereby maximizing the aggregated data throughput of all user nodes within a limited flight time. Furthermore,[9269485] shows that, by constructing a signal-to-interference-plus-noise ratio (SINR) map, the Dijkstra algorithm can be employed to identify the shortest path that consistently satisfies communication quality requirements while avoiding co-channel interference from neighboring terrestrial BS. In addition, the authors of[9354009] employed deep reinforcement learning to plan the trajectory of cellular-connected UAVs, and simultaneously constructing CKM to improve the training efficiency of the learning agent. Meanwhile, existing studies have demonstrated that employing CKM in BS placement or configuration can yield notable performance gains and cost reductions. For example,[11063460] presents a CGM-based scheme that optimizes the antenna downtilt and azimuth angles to maximize cell coverage and [10008712] proposes a CGM-enabled BS deployment method that satisfies multi-BS coverage requirements and quality-of-service constraints. Since CKMs can accurately characterize the spatial properties of wireless channels, their applications can be extended to higher-layer network functionalities such as user resource allocation and scheduling. The authors in[9104036] develop a neural-network–based power control scheme leveraging large-scale fading coefficients. Hence, constructing multi-cell CGMs can provide valuable support for resource allocation and power control in multi-cell massive MIMO systems. Moreover, as shown in[10474197], CKM can be employed in cell-free massive MIMO (CF-mMIMO) systems to facilitate access-point (AP) switching, thereby mitigating pilot contamination and reducing energy consumption associated with path-loss estimation. Beyond the aforementioned applications, CKM has also attracted increasing attention in a broader range of areas, such as ISAC[10693799], physical layer security[8815467], intelligent beamforming[10287775], and device-to-device communications[9053347].
@@
Considering that obtaining complete channel information over an entire spatial region purely through measurements is costly, time-consuming, and practically infeasible, research efforts have been devoted to developing efficient CKM construction methods. Depending on the implementation approach, existing CKM construction techniques can be categorized into four types: model-based methods using statistical channel models, model-based methods using deterministic channel models, spatial interpolation–based methods, and deep learning–based methods. The statistical channel model-based construction methods refer to utilizing existing statistical channel models to compute large-scale or small-scale fading channel information. Commonly used statistical channel models include the log-distance model[5635444] and the Okumura–Hata model[1622772] for large-scale fading computation, as well as Quadriga[6758357] and the 6G PCM (pervasive channel model)[10225614] for small-scale fading modeling. Although this method is simple, it often yields CKMs with limited accuracy. The deterministic channel model–based construction methods obtain channel-related information by performing ray-tracing–based channel simulations. Representative ray-tracing simulators include Wireless InSite[RemcomInc2024], WinProp[7916282], and Sionna[Hoydis2022SionnaAO], among others. Although this approach can achieve high accuracy, it suffers from extremely high computational complexity, making it difficult to meet the real-time requirements of physical-layer algorithms.
@@
The spatial interpolation–based approach constructs CKMs using a limited number of measured channel samples at selected locations. By exploiting the spatial correlation, smoothness, or low-rank characteristics of the wireless channel over the spatial domain, classical interpolation algorithms are applied to estimate the channel information in unmeasured regions. Typical spatial interpolation algorithms used for CKM construction include inverse distance weighting (IDW)[Lu2008], Kriging[1371308], radial basis function (RBF) interpolation[Bishop1995], matrix completion[7472907], compressed sensing[5352337], and the expectation–maximization (EM) algorithm[9771802]. Furthermore, the number of samples required to construct a CKM using spatial interpolation methods has been theoretically analyzed in [10530520]. Although the above methods significantly reduce the computational complexity compared with ray-tracing–based approaches, they do not incorporate physical environmental information. Moreover, these spatial interpolation algorithms rely on the assumptions of strong spatial correlation, smoothness, or low-rank characteristics of the channel. In densely built urban areas, where such assumptions are often violated, the accuracy of CKMs obtained through interpolation can degrade considerably.
@@
Deep learning models can effectively establish a mapping from the physical environment map to CKM, thereby enabling more accurate and generalizable CKM construction. The earliest attempts to apply deep learning for CKM construction were inspired by models in the field of computer vision. Specifically, convolutional neural networks based on autoencoder architectures [10682525, 9523765], UNet structures[9354041], and Laplacian pyramid frameworks[10791446] were introduced to capture the spatial correlations of large-scale fading channels.
@@
The discriminative models employed in the aforementioned studies typically yield over-smoothed predictions that lack sufficient high-frequency details. However, the spatial distribution of channel-related features in complex environments exhibits strong nonlinearity and fine-grained variations, making it difficult for purely supervised learning to accurately capture such characteristics. In contrast, generative models learn the underlying data distribution and are therefore capable of capturing higher-order statistical characteristics, enabling more accurate prediction of fine-grained spatial variations in channel information. Previous studies[8794603, 10130091] have developed conditional GAN (cGAN)–based frameworks to achieve environment-consistent CKM generation. However, the GAN framework involves a minimax optimization objective, which often leads to unstable training and convergence difficulties. In recent years, diffusion models have been recognized as a more powerful class of generative models, providing significantly higher generation quality than GANs. Inspired by their success in other domains, researchers have introduced conditional diffusion frameworks for CKM construction. The studies in[fu2025ckmdiff, 10764739] have demonstrated that conditional diffusion models can achieve state-of-the-art CKM prediction accuracy under various input conditions.
@@
However, the aforementioned studies that employ diffusion models for CKM construction directly transfer conditional diffusion frameworks from the field of computer vision. In image generation tasks, various improved conditional diffusion models are designed to produce images that are more realistic and visually appealing. Consequently, their performance is typically evaluated using image quality metrics such as the structural similarity index measure(SSIM), Fréchet inception distance(FID), and learned perceptual image patch similarity. In contrast, the primary objective in CKM construction is to achieve high prediction accuracy rather than visual fidelity. Therefore, conditional diffusion models with high image generation quality are not suitable for direct application to CKM construction. In this paper, motivated by the physical characteristics of wireless signal propagation, we integrate physical prior knowledge with the diffusion model and propose a physics-inspired diffusion-based CKM generation framework. The proposed model not only inherits the advantages of diffusion models in generating high-quality results with richer high-frequency details, but also ensures that the generated CKMs conform to the physical constraints of wireless signal propagation. It is worth noting that several physics-informed methods for constructing CGM have been proposed in[10608081] and[shahid2025reveal]. These approaches introduce physical constraints derived from Maxwell’s equations and incorporate them into the neural network training process as additional loss terms. However, we note that Maxwell’s equations represent the microscopic formulation of wireless signal propagation, as the described electric field variations occur at the wavelength scale. Therefore, they are not well-suited to serve as physical constraints for large-scale fading, which characterizes the macroscopic propagation behavior of wireless signals. Moreover, the boundary conditions of electromagnetic reflections, refractions, and diffractions are highly complex, and existing physics-inspired methods do not impose boundary-condition residuals as training constraints, further reducing the physical fidelity of the CGM construction. To address these limitations, this paper heuristically designs three novel physical constraint terms that better characterize the spatial distribution of large-scale fading, and innovatively integrates them into the diffusion model framework.
@@
Based on the above discussions, the main contributions of this paper can be summarized as follows:
@@
    To address the insufficient physical consistency exhibited by existing CGM construction methods, we design three physics-inspired regularization terms, namely the edge loss, the regional propagation loss, and the multi-scale feature loss. These three regularizations respectively reflect distinct physical phenomena exhibited by electromagnetic waves when encountering obstacles, and are therefore closely tied to the propagation environment. This makes them particularly suitable for CGM construction in complex scenarios such as urban canyons and indoor environments.
@@
2 CKM Construction Problem Description
The objective of this paper is to construct a classical type of CKM—the CGM. In the construction of CKM, including the CGM, the map is typically divided into $H\times H$ grids. Consequently, both the ground-truth CGM $\mathbf{x}_{GT}$ and the predicted CGM $\mathbf{x}_{pred}$ are represented as real-valued matrices of size $H\times H$. Accordingly, when external conditional information is required for CGM construction, the input data are also represented as real-valued matrices of the same size. In the transmitter location matrix $\mathbf{L}_{\text{TX}}\in\mathbb{R}^{H\times H}$, only the element corresponding to the transmitter position is set to 1, while all other elements are 0. In the building environment matrix $\mathbf{B}\in\mathbb{R}^{H\times H}$, the elements representing building regions are set to 1, and all others are 0. In the vehicle location matrix $\mathbf{C}\in\mathbb{R}^{H\times H}$, the element corresponding to the vehicle position is set to 1, with all remaining elements being 0. In the sparse observation matrix $\mathbf{S}\in\mathbb{R}^{H\times H}$, the elements at the observation positions contain the actual observed values, whereas all other elements are set to 0.
@@ 2 CKM Construction Problem Description
3 Theoretical Foundations of Diffusion Models
@@
3.1 Denoising Diffusion Probabilistic Model
Inspired by nonequilibrium thermodynamics, Denoising Diffusion Probabilistic Model (DDPM)[ho2020denoising] formulates the forward noise-adding and reverse denoising processes as Markov chains. Therefore, DDPM mainly consists of two components: a forward process that gradually adds noise to the input, and a reverse process that learns how to denoise it.
@@ 3.1 Denoising Diffusion Probabilistic Model
$q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{x}_{0})$ $=\frac{q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1},\mathbf{x}_{0})q(\mathbf{x}_{t-1}\mid\mathbf{x}_{0})}{q(\mathbf{x}_{t}\mid\mathbf{x}_{0})}\propto\mathcal{N}\left(\mathbf{x}_{t-1};\tilde{\mu}_{t},\tilde{\beta}_{t}\mathbf{I}\right),$ (6)
@@
3.2 Denoising Diffusion Implicit Model
The DDPM requires $T$ iterative steps to generate a single sample, which results in relatively slow generation speed. To address this limitation, the Denoising Diffusion Implicit Model (DDIM)[Song2020DenoisingDI] removes the constraint that the reverse denoising process must follow a Markov chain, thereby achieving significantly faster sampling while maintaining comparable generation quality.